Evaluate the definite integral. Enter exact values. 2 cos (2) dc V1 - 22 - =
Answer #1

I am attaching the pictures of solution. If you find my answer helpful please press thumbs up button. Thanks

Ans let I- s - 2 105(x) da la x² Now let cos/²) = -6 > x= cos (4) 3Now by del (105\/3)] -4,4) = dt da Jox (became en tot mom) - 1-bey = da= at ہوا ST-xdso 2005(x) da 5l-x - 2 Cost(x) 2.fo lay ر- اکا -2 S + dt (using , ©) -2 + (as stat t 2 + m-4 t t = (ost(x)) So (Cost (x)) (bu={cos (137 - (cos (-1}} (0)2- (O) 2 as os (os (1) = 0 COST(-1) = 1 F- sosten 2 Cos (0) rl-x teral -1